D5 Nine teenfortyfour to Nineteensixtynine Nineteenseventy to Twentythirteen 10 12 13 14 16 17 18 19 20 21 0 24 25 26 27 28 29 30 31 32 0 34 35 36
301 32 12110561353 32 36722525420 2260 251 3 1 0 3 s- 9 0 1 2 3 4 5 6 7 8 9 0 1 2 3 4 5 6 7 8 9 0 1 2 3 4 5 46 47 7
1. There is a csv file named Hurricanes. Input that file into R. 2. Create a variable for the hurricanes between 1944 and 1960 and another variable for the hurricanes between 1970 and 2013 Create two boxplots from the data. Do the boxplots indicate that the means are the same or different? State the null and alternative hypothesis: Use R to test your hypothesis What is your conclusion. 3. 4. 5. 6.
Show R work and conclusions thanks.
Answer #1

#1. & 2.

> #variable A for 1944-1960
> #variable B for 1970-2013
> A=c(3,2,1,2,4,3,7,2,3,3,2,5,2,2,4,2,2,6,0,2,5,1,3,1,0,3)
> B=c(2,1,0,1,2,3,2,1,2,2,2,3,1,1,1,3,0,1,3,2,1,2,1,1,0,5,6,1,3,5,3,3,2,3,6,7,2,2,5,2,5,4,2,0)
#3.

> par(mfrow=c(1,2))
> boxplot(A,main="A")
> boxplot(B,main="B")
xqponMFg+H5AAAAAElFTkSuQmCC

#BOXPLOTS indicate that means are same.

#4.

#Null hypothesis \mu _{A}=\mu _{B}

#Alternative hypothesis \mu _{A}\neq \mu _{B}

#5.

> var.test(A,B)

F test to compare two variances

data: A and B
F = 0.97515, num df = 25, denom df = 43, p-value = 0.9692
alternative hypothesis: true ratio of variances is not equal to 1
95 percent confidence interval:
0.4955629 2.0523066
sample estimates:
ratio of variances
0.9751496

#p-value=0.9692 which is greater than 0.05. Therefore, variances of A and B are homogeneous.

#Student's two sample t-test is to be carried out.

> t.test(A,B,var.equal=TRUE)

Two Sample t-test

data: A and B
t = 0.77921, df = 68, p-value = 0.4386
alternative hypothesis: true difference in means is not equal to 0
95 percent confidence interval:
-0.5130186 1.1703613
sample estimates:
mean of x mean of y
2.692308 2.363636

#p-value=0.4386 which is greater than 0.05. Therefore, means of A and B are equal.

#6.

#Conclusion: Null Hypothesis is accepted. Means for A (1944-1960) and B (1970-2013) are equal.